OpenStudy (anonymous):

If f is the solution of x*f'(x)-f(x)=x such that f(-1)=1, find f(e^-1). Answer: -2e^-1.

OpenStudy (anonymous):

damn lost all my typing :((

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i got y = x ln(x) - x f(e^-1) = (e^-1) ln(e^-1) - e^-1 = -e^-1 - e^-1 = -2e^-1

OpenStudy (anonymous):

first order linear differential equation.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no, you can't just solve for y'

OpenStudy (anonymous):

that's not how you solve first order linear differential equation

OpenStudy (anonymous):

well your answer didn't match with the given answer

OpenStudy (anonymous):

ok, y' - (1/x) y = 1

OpenStudy (anonymous):

multiply both sides by 1/x to reverse product rule (1/x)y' - (1/x^2)y = 1/x d/dx (1/x)y = 1/x (1/x)y = ln|x| + C y = x ln|x| + Cx given f(-1) = 1 1 = (-1) ln|-1| + C(-1) 1 = -C C = -1 y = x ln|x| - x plug in e^-1 for x gives -2e^-1

OpenStudy (anonymous):

have you taken differential equation before?

OpenStudy (anonymous):

Yes. But this problem isn't Differential Equations.